can you please assist me with the following problems
1. use elimination to solved
x/3+4y=30
2x-y=5
6x-y/4=-9
3x+2y=-30
use subsitution to solved
y=x2+2x-2
y=3x+4
Found 2 solutions by pramodkumar, malakumar_kos@yahoo.com: Answer by pramodkumar(2) 
 1)
x/3+4y=30 ---- (i)
2x-y=5 ---- (ii)
from (i)X6 - (ii)
2x+24y= 180
2x-y = 5
- + -
------------
25y = 175
or, y = 7
substitution the value of x in (ii)
2x-7 = 5
or, 2x = 12
or, x = 6
Ans. x = 6
y = 7
________________________________________________
6x-y/4=-9 ---- (i)
3x+2y=-30 ---- (ii)
(i)X8 + (ii) we get
48x-2y = -72
3x+2y = -30
----------------
51x = -102
or x = -2
put the value of x in (ii)
-6+2y = -2
or 2y = -2+6
or y = 2
Ans. x = -2, y = 2
2)
y=x2+2x-2 --- (i)
y=3x+4 --- (ii)
substitution the value of y from (ii) in (i), we get
3x+4 = x2+2x-2
or, x2-x-6=0
or, x2-3x+2x-6=0
or, (x-3)(x+2)=0
i.e. x=3 or x=-2
from (ii)
y = 13 or y = -2
Answer by malakumar_kos@yahoo.com(315) 
 1)x/3+4y = 30......eq'n(1)
2x-y = 5 ........eq'n(2)
multiply eq'n(1) by 3 throughout we get x+12y = 90 hence x = 90-12y
substitute for x in eq'n (2) we get 2(90-12y)-y = 5
180- 24y-y = 5
180-5 = 25y
25y = 175 or y = 175/25
y = 7
substitute for y in 2x-y = 5
2x = 5+y = 5+7 = 12 i.e,x = 12/2 = 6
the solution for the eq'ns are x = 6 and y = 7
2)6x-y/4 = -9........eq'n (1)
3x+2y = -30........eq'n(2)
multiply eq'n (1) by 4 throughout we get 24x-y = -36
24x+36 = y
substitute for y in eq'n(2) we get 3x+2(24x+36) = -30
3x+48x+72= -30
51x = -30-72 = -102
x = -102/51 = -2
substituting for x in 3x+2y =-30 we get 3(-2)+2y = -30
-6 +2y -30 or 2y = -30+6 = -24
y = -24/2 = -12
the solution for the eq'ns is x = -2 and y = -12
3)y = x^2+2x-2.........eq'n (1)
y = 3x+4 ............eq'n(2)
substituting the value of y from eq'n(2) in eq'n(1)
we get 3x+4 = x^2+2x-2 or x^2+2x-3x-2-4 = 0
x^2-x-6 = 0 spilliting -6 = -3*+2
x^2-3x+2x-6=0
x(x-3)+2(x-3) = 0
(x-3).(x+2) = 0 or x-3 = 0, x=3
x+2 = 0 , x = -2
if x=-2then value of y = 3(-2)+4 = -6+4 = -2
if x = 3 then the value of y = 3(3)+4 = 9+4 = 13
| |
|